OpenStudy (anonymous):

find the zeros of each function. 1. f(x)= x^2-7x 2. f(x)= x^2 -9x+20

OpenStudy (anonymous):

plsssss find the zeros of each function. 1. f(x)= x^2-7x 2. f(x)= x^2 -9x+20

OpenStudy (hauntedwoodsgal):

I think I got it!!!!!